OpenStudy (anonymous):

Find the derivative of F(X)= 2X^2 -16x+35.. I know how to plug in the f'(X) equation but I get stuck at: 2X^2 + 4xh+2h^2 -16x-16h+35-2x^2 16x-35.... How did they get the 4xh. I understand how they got the 2x^2 and the 2h^2 but how does 2(x+h)^2 give u a 4hx?

OpenStudy (loser66):

so, tell me 2( a+b)^2 = 2 (a^2 +2ab + b^2) = 2a^2 + 4ab + 2b^2. yours is x and h, not a, b as mine, but apply the same rule, that's why you have 4xh